surface Area and volume of spheres, cones and cylinders
16. The diameter of a volleyball is 10.9cm
a) Determine the amount of leather needed to make the ball
b)Determine the amount of air inside the ball

Answer by stanbon(75887)

The diameter of a volleyball is 10.9cm
a) Determine the amount of leather needed to make the ball
Area = 4(pi)r^2
Area = 4(pi)(5.45)^2 = 373.25 sq cm.
------------------------------------------
b)Determine the amount of air inside the ball
Volume = (4/3)(pi)r^3
Volume = (4/3)(pi)(5.45)^3 = 678.08 cm^3
=============================================
Cheers,
Stan H.
=======